KalinaCriss666:

A box of pencils costs $10.The total cost (c) as a function of the number of boxes (b) is expressed as c = f(b) = 10b. What is the domain of this function? A} all positive integers and zero B} all real numbers C} all real numbers except 0 D} all positive real numbers except 10

Shadow:

Do we have answer choices?

Shadow:

There we go, haha

Shadow:

@KalinaCriss666 Do you know what domain is?

KalinaCriss666:

Nope, I completely forgot

Shadow:

In a classic slope equation: \[y = mx + b\] \[f(x) = mx + b\] The domain essentially represents all of the possible x values. So in \[c = f(b) = 10b\] The domain represents all of the possible values for b.

KalinaCriss666:

OMFL THANKS SO MUCH!!!

Shadow:

You got it?

KalinaCriss666:

Yes!!

Shadow:

What answer did you get?

Shadow:

@KalinaCriss666

KalinaCriss666:

Eh, sorry, gimme a minute

KalinaCriss666:

I missed four questions, so I'm wondering which one I actually got right XD

Shadow:

Well the answer to this one was A, lol

KalinaCriss666:

Well f, I chose D X'D

Shadow:

D: Why

KalinaCriss666:

My brain is f'ed up today, my dude

Shadow:

Think of it this way. All real numbers means both negative and positive numbers. Can there ever be -5 boxes of pencils?

Shadow:

Remember what b stands for, # of boxes

Shadow:

There can however, be 0 boxes and 3 boxes. Thus why it is positive integers and zero. Also, integers basically mean it's not a fraction. So no 1/2 boxes or 3.5 boxes.